Using traditional methods, it takes 8.78.7 hours to receive a basic flying license. A new license training method using Computer Aided Instruction (CAI) has been proposed. A researcher used the technique with 1212 students and observed that they had a mean of 8.58.5 hours with a standard deviation of 1.11.1. A level of significance of 0.050.05 will be used to determine if the technique performs differently than the traditional method. Assume the population distribution is approximately normal. Find the value of the test statistic. Round your answer to three decimal places.
Answer #1
population mean μ= 8.7
sample mean 'x̄= 8.500
sample size   n= 12.00
sample std deviation s= 1.100
std error 'sx=s/√n= 0.318
test statistics t ='(x-μ)*√n/sx= -0.630
p value      = 0.5417

(since p value >0.05 , we fail to reject the null hypothesis()